OpenStudy (heart_offire):

Divide. 16x3 - 8x2 + 4x4 ------------ 2x A) 8x^2 - 4x + 2x^3 B) 8x^2 - 4 + 2x^3 C) 8x^3 - 4x + 2x^3 D) 8x^2 - 4x + 2x

OpenStudy (neonumbrella5115):

Have you learned the synthetic division or long division method?

OpenStudy (heart_offire):

well i learned how to just break it apart in three since there is three numbers on top and one at the bottom so it would be 16x^3/2x etc.

OpenStudy (tkhunny):

Super. Then do that.

OpenStudy (neonumbrella5115):

Interesting. I didn't learn it that way however I think I can help you do specifically this problem the way you want. So when you divide the expression you said in the post above, you get 8x^2, but that doesn't help you narrow down your choices. Right?

OpenStudy (heart_offire):

well i did that but it got confusing.......right @neonumbrella5115

OpenStudy (tkhunny):

Demonstrate.

OpenStudy (heart_offire):

my method doesnt always work with every equation like this one

OpenStudy (neonumbrella5115):

So next, you would do -8x^2/ 2x and you would get: @heart_offire

OpenStudy (heart_offire):

4x^2

OpenStudy (neonumbrella5115):

Almost, since you are dividing x^2 by an x there would only be one x in your answer.

OpenStudy (heart_offire):

oh.. thought they stay the same since its division

OpenStudy (neonumbrella5115):

That's a mistake many people make, so no worries! When you're working with dividing exponents you subtract the exponent #. Example x^6/x^3 is x^3 So you want to try the last one on your own?

OpenStudy (heart_offire):

2x^3

OpenStudy (heart_offire):

4x^4/2x=2x^3

OpenStudy (tkhunny):

Division always works. \(\dfrac{16x^{3} - 8x^{2} + 4x^{4}}{2x} = \dfrac{2x(8x^{2} - 4x +2x^{3})}{2x} = \dfrac{2x}{2x}(8x^{2} - 4x +2x^{3})\)

OpenStudy (neonumbrella5115):

Yup! So the only choice with 4x and 2x^3 and 8x^2 is...

OpenStudy (heart_offire):

A

OpenStudy (heart_offire):

that was my first original chioce but i clicked C instead i have no idea why though

OpenStudy (neonumbrella5115):

@tkhunny That's how I usually solve these kinds of problems, as it always works no matter what. It also is very simple, once you get the hang of it. If you continue to do problems like these @heart_offire I would suggest learning the divison method above^^

OpenStudy (neonumbrella5115):

A is correct!! Good job! If you need anymore help just tag me!

OpenStudy (heart_offire):

^sure i'll tag you with the next problem im having trouble solving